OpenStudy (anonymous):

Let log a=x, log b=y,and log c=z. Express log (a^2b)/c^3 in terms of x,y,and z.

OpenStudy (anonymous):

Do you mean \(log (\frac{a^2b}{c^3})\) for your question?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

log (ab) = log a + log b log (a/b) = log a - log b log (a^n) = n log a So, \[\log \frac{a^2b}{c^3}\]\[=\log (a^2b) - \log c^3\]\[=\log a^2+ \log b - \log c^3\]\[=2\log a+ \log b - 3\log c\]\[=...?\]

OpenStudy (anonymous):

2x+y-3z ?

OpenStudy (anonymous):

Is that correct?